# Bifröst: 3D-Aware Image compositing with Language Instructions

### v1

@@ Abstract
This paper introduces Bifröst, a novel 3D-aware framework that is built upon diffusion models to perform instruction-based image composition. Previous methods concentrate on image compositing at the 2D level, which fall short in handling complex spatial relationships (e.g., occlusion). Bifröst addresses these issues by training MLLM as a 2.5D location predictor and integrating depth maps as an extra condition during the generation process to bridge the gap between 2D and 3D, which enhances spatial comprehension and supports sophisticated spatial interactions. Our method begins by fine-tuning MLLM with a custom counterfactual dataset to predict 2.5D object locations in complex backgrounds from language instructions. Then, the image-compositing model is uniquely designed to process multiple types of input features, enabling it to perform high-fidelity image compositions that consider occlusion, depth blur, and image harmonization. Extensive qualitative and quantitative evaluations demonstrate that Bifröst significantly outperforms existing methods, providing a robust solution for generating realistically composed images in scenarios demanding intricate spatial understanding. This work not only pushes the boundaries of generative image compositing but also reduces reliance on expensive annotated datasets by effectively utilizing existing resources in innovative ways.
@@ 1 Introduction
Image generation has flourished alongside the advancement of diffusion models (Song et al., 2021; Ho et al., 2020; Rombach et al., 2022; Ramesh et al., 2022). Recent works (Saharia et al., 2022; Liu et al., 2023b; Brooks et al., 2023; Zhang et al., 2023b; Huang et al., 2023; Li et al., 2023b; Chen et al., 2024) add conditional controls, e.g., text prompts, scribbles, skeleton maps to the diffusion models, offering significant potentials for controllable image editing. Among these methods for image editing, generative object-level image compositing (Yang et al., 2023; Song et al., 2023; Chen et al., 2024; Song et al., 2024) is a novel yet challenging task that aims to seamlessly inject an outside reference object into a given background image with a specific location, creating a cohesive and realistic image. This ability is significantly required in practical applications including E-commerce, effect-image rendering, poster-making, professional editing, etc.
Achieving arbitrary personalized object-level image compositing necessitates a deep understanding of the visual concept inherent to both the identity of the reference object and spatial relations of the background image. To date, this task has not been well addressed. Paint-by-Example (Yang et al., 2023) and Objectstitch (Song et al., 2023) use a target image as the template to edit a specific region of the background image, but they could not generate ID-consistent contents, especially for untrained categories. On the other hand, (Chen et al., 2024; Song et al., 2024) generate objects with ID (identity) preserved in the target scene, but they fall short in processing complicated 3D geometry relations (e.g., the occlusion) as they only consider 2D-level composition. To sum up, previous methods mainly either 1) fail to achieve both ID preservation and background harmony, or 2) do not explicitly take into account the geometry behind the background and fail to accurately compose objects and backgrounds in complex spatial relations.
@@
We divide the training procedure into two stages. In the first stage, we finetune an MLLM (e.g., LLaVA (Liu et al., 2023a, 2024)) for 2.5D predictions of objects in complex scenes with language instructions. In the second stage, we train the image composition model. To compose images with complicated spatial relationships, we introduce depth maps as conditions for image generation. In addition, we leverage an ID extractor to generate discriminative ID tokens and a frequency-aware detail extractor to extract the high-frequency 2D shape information (detail maps) for ID preservation. We unify the depth maps, ID tokens, and detail maps to guide a pre-trained text-to-image diffusion model to generate desired image composition results. This two-stage training paradigm allows us to utilize a number of existing 2D datasets for common visual tasks e.g., visual segmentation, and detection, avoiding collecting large-volume text-image data specially designed for arbitrary object-level image compositing.
@@ 2.1 Image compositing with Diffusion Models
Image compositing, an essential technique in image editing, seamlessly integrates a reference object into a background image, aiming for realism and high fidelity. Traditional methods, such as image harmonization (Jiang et al., 2021; Xue et al., 2022; Guerreiro et al., 2023; Ke et al., 2022) and blending (Pérez et al., 2003; Zhang et al., 2020, 2021; Wu et al., 2019) primarily ensure color and lighting consistency but inadequately address geometric discrepancies. The introduction of diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015; Song et al., 2021; Rombach et al., 2022) has shifted focus towards comprehensive frameworks that address all facets of image compositing. Methods like (Yang et al., 2023; Song et al., 2023) often use CLIP-based adapters to utilize pretrained models, yet they compromise the object’s identity preservation, focusing mainly on high-level semantic representations.More recent studies prioritize maintaining the appearance in generative object compositing. Notable developments in this field include AnyDoor (Chen et al., 2024) and ControlCom (Zhang et al., 2023a). AnyDoor integrates DINOv2 (Oquab et al., 2023) with a high-frequency filter, while ControlCom introduces a local enhancement module, both improving appearance retention. However, these approaches still face challenges in spatial correction capabilities. Most recent work IMPRINT (Song et al., 2024) trains an ID-preserving encoder that enhances the visual consistency of the object while maintaining geometry and color harmonization. However, none of the work can deal with composition with occlusion and more complex spatial relations.In contrast, our models novelly propose a spatial aware generative model that allows more accuracy and complex image compositing while maintaining geometry and color harmonization.
@@ 3 Method
Figure 5: Overview of training pipeline of Bifröst on image compositing stage. A segmentation module is first adopted to get the masked image and object without background, followed by an ID extractor to obtain its identity information. The high-frequency filter is then applied to extract the detail of the object, stitch the result with the scene at the predicted location, and employ a detail extractor to complement the ID extractor with texture details. We then use a depth predictor to estimate the depth of the image and apply a depth extractor to capture the spatial information of the scene. Finally, the ID tokens, detail maps, and depth maps are injected into a pre-trained diffusion model to produce the final synthesis, where the target object favorably blends with its surroundings yet satisfies complex spatial relations.
@@
where $\mathcal{E}_{d}$ is the depth extractor and $\boldsymbol{c}_{\mathrm{d}}$ is the extracted depth condition. In our experiment, the detail and depth extractors are ControlNet-style UNet encoders, producing a series of detail maps with hierarchical resolutions.
@@
$\left.\mathcal{L}_{compose}=\mathbb{E}_{\boldsymbol{z}_{0},\boldsymbol{t},\boldsymbol{c}_{i},\boldsymbol{c}_{\mathrm{f}},\epsilon\sim\mathcal{N}(0,1)}\left[\|\epsilon-\epsilon_{\theta}\left(\boldsymbol{z}_{t},\boldsymbol{t},\boldsymbol{c}_{i},\boldsymbol{c}_{\mathrm{f}}\right)\right)\|_{2}^{2}\right],$ (5)
where $\boldsymbol{c}_{f}=\boldsymbol{c}_{h}+\lambda\cdot\boldsymbol{c}_{d}$ and $\lambda$ is a hyper parameter weight between two controls. Specifically, the ID tokens are injected into each UNet layer via cross-attention. For the high-frequency detail maps and depth maps, we add them together first and concatenate them with UNet decoder features at each resolution. During training, we freeze the pre-trained parameters of the UNet encoder to preserve the priors and tune the UNet decoder to adapt it to our new task.
@@
Dataset Generation. The ideal training samples are image pairs for “the same object in different scenes with different poses”, which existing datasets do not provide. Previous works (Yang et al., 2023; Song et al., 2023) use single images with augmentations like rotation and flip, which are insufficient for realistic pose and view variants. To address this, we use video datasets to capture frames of the same object as complementary following (Chen et al., 2024; Song et al., 2024). The data preparation pipeline is demonstrated in Fig. 6. For a video, we pick two frames and extract the masks for the foreground object. Then, we mask the background for one image and crop it around the mask as the reference object. For the other frame, we use the masked image with an augmented mask shape as the background image, and the unmasked image serves as the training ground truth. The dataset quality is another key to better identity preservation and pose variation. The full data used is listed in Tab. 1, which covers a large variety of domains such as nature scenes (SAM (Kirillov et al., 2023), LVIS (Gupta et al., 2019), HFlickr (Cong et al., 2020), DUT (Wang et al., 2017), and MSRA-10K (Borji et al., 2015)), panoptic video segmentation datasets (YoutubeVOS (Xu et al., 2018), VIPSeg (Miao et al., 2022), and MOSE (Ding et al., 2023)), and virtual try-on dataset (VitonHD (Choi et al., 2021)).
Figure 6: Data preparation pipeline of leveraging videos. Given a clip, we first sample two frames and select one instance from one frame as the reference object, and treat the same instance on the other frame as the training supervision.
@@ 4 Experiment
Table 4: User study on the comparison between our Bifröst and existing alternatives. “Quality”, “Fidelity”, and “Diversity” measure synthesis quality, object identity preservation, and object local variation (i.e., across four proposals), respectively. Each metric is rated from 1 (worst) to 5 (best).
@@ 4.1 Implementation Details
Zoom-in strategy. During inference, given a scene image and a location box, we expand the box into a square with an amplifier ratio of 2.0. Then, we crop the square and resize it to $512\times 512$ as the input for our diffusion model. Thus, we could deal with scene images with arbitrary aspect ratios and boxes for extremely small or large areas.
@@
Evaluation metrics. We test the IoU and MSE loss to evaluate the accuracy of the predicted bounding box and depth of our MLLM. For the image compositing model, on our constructed DreamBooth dataset, we follow DreamBooth (Ruiz et al., 2023) to calculate the CLIP-Score and DINO-Score, as these metrics could reflect the similarity between the generated region and the target object. We also calculate the FID (Heusel et al., 2017) to evaluate the realism, which indicates the compositing quality. In addition, we organize user studies with a group of 30 annotators to rate the generated results from the perspective of fidelity, quality, diversity, and 3D awareness.
@@ 5 Conclusion
Acknowledgements This work is partially supported by the National Natural Science Foundation of China (Grant No. 8326014), CUHK Direct Grants (Grant No. 4055190), and The Shun Hing Institute of Advanced Engineering (SHIAE) No. 8115074. We thank Boqing Gong (Google) for feedback on an early draft of this paper and Yuanyuan Zhang (University of Liverpool) for proposing the name Bifröst of our model.
@@ Appendix A Implementation Details
For spatial aware image composing model. We use Stable Diffusion 2.1 as the pre-trained diffusion model. The model can also be initialized from the pre-trained weights of  (Chen et al., 2024) to better utilize the prior knowledge. We choose Adam (Kingma and Ba, 2014) optimizer with an initial learning rate of $1e^{-5}$. The DINO-V2 (Oquab et al., 2023) ID extractor and the encoder of the U-net are frozen during the training. The decoder of the U-net, detail extractor, and depth extractor are fine-tuned during this process. The batch size is set as 16, we train 20k steps on 4 A800 GPUs, which takes 1 day to finish the training.
@@ B.1 MLLM Fine Tuning
We leverage the pre-trained visual encoders to extract the identity information of the target object. Instead of using CLIP (Radford et al., 2021) image encoder to embed the target object (Yang et al., 2023; Song et al., 2023), previous works (Chen et al., 2024; Song et al., 2024) has shown that DINO-V2 (Oquab et al., 2023) has strong ability to preserve more discriminative features and project the object into an augmentation-invariant feature space. Thus, we first use pre-trained DINO-V2 to encode image as a global token $\mathbf{T}_{\text{g}}^{1\times 1536}$, and patch tokens $\mathbf{T}_{\text{p}}^{256\times 1536}$. We concatenate the two types of tokens to preserve more information. Following (Chen et al., 2024), we use a single linear layer to bridge embedding space of DINOV2 and the pre-trained text-to-image UNet. The final projected ID tokens are $\mathbf{T}_{\text{ID}}^{257\times 1024}$.
Collage representation. Unlike previous works (Casanova et al., 2023; Sarukkai et al., 2024; Chen et al., 2024) that either use the image or high-frequency map with colors as the complementary information. We found that the generated objects are similar to the reference image and can not adapt the color and light condition to fuse into the background seamlessly which results in a “copy-past” style. Specifically, we remove the color information of the HF-map in (Chen et al., 2024) to represent fine-grained details of the object, which yields visually more coherent and natural results.
Mask Shape Augmentation. As discussed in Sec. 3.2, to enable more user control, we define five levels of coarse masks, including the bounding box mask. Fig. 14 shows all the mask types. As the coarse level increases (from mask 1 to mask 5), the model has more freedom to generate the object.
@@
The total process of Detail Extract can be formulated as
@@
where $\mathbf{K}_{h},\mathbf{K}_{v}$ denote horizontal and vertical Sobel (Kanopoulos et al., 1988) kernels, acting as high-pass filters. $\otimes$,$\odot$ , ⊙ refer to convolution and Hadamard product. Given an object image $\mathbf{I}_{obj}$, we first extract the high-frequency regions using these high-pass filters and add a shape-augmented mask $\mathbf{M}_{\text{aug}}$ to filter out the information near the outer contour of the target object.

### v2

@@ Abstract
This paper introduces Bifröst, a novel 3D-aware framework that is built upon diffusion models to perform instruction-based image composition. Previous methods concentrate on image compositing at the 2D level, which fall short in handling complex spatial relationships (e.g., occlusion). Bifröst addresses these issues by training MLLM as a 2.5D location predictor and integrating depth maps as an extra condition during the generation process to bridge the gap between 2D and 3D, which enhances spatial comprehension and supports sophisticated spatial interactions. Our method begins by fine-tuning MLLM with a custom counterfactual dataset to predict 2.5D object locations in complex backgrounds from language instructions. Then, the image-compositing model is uniquely designed to process multiple types of input features, enabling it to perform high-fidelity image compositions that consider occlusion, depth blur, and image harmonization. Extensive qualitative and quantitative evaluations demonstrate that Bifröst significantly outperforms existing methods, providing a robust solution for generating realistically composited images in scenarios demanding intricate spatial understanding. This work not only pushes the boundaries of generative image compositing but also reduces reliance on expensive annotated datasets by effectively utilizing existing resources in innovative ways.
@@ 1 Introduction
Image generation has flourished alongside the advancement of diffusion models (Song et al., 2021; Ho et al., 2020; Rombach et al., 2022; Ramesh et al., 2022). Recent works (Saharia et al., 2022; Liu et al., 2023b; Brooks et al., 2023; Zhang et al., 2023b; Huang et al., 2023; Li et al., 2023; Chen et al., 2024; He et al., 2024) add conditional controls, e.g., text prompts, scribbles, skeleton maps to the diffusion models, offering significant potentials for controllable image editing. Among these methods for image editing, generative object-level image compositing (Yang et al., 2023; Song et al., 2023; Chen et al., 2024; Song et al., 2024) is a novel yet challenging task that aims to seamlessly inject an outside reference object into a given background image with a specific location, creating a cohesive and realistic image. This ability is significantly required in practical applications including E-commerce, effect-image rendering, poster-making, professional editing, etc.
Achieving arbitrary personalized object-level image compositing necessitates a deep understanding of the visual concept inherent to both the identity of the reference object and spatial relations of the background image. To date, this task has not been well addressed. Paint-by-Example (Yang et al., 2023) and Objectstitch (Song et al., 2023) use a target image as the template to edit a specific region of the background image, but they could not generate ID-consistent contents, especially for untrained categories. On the other hand, (Chen et al., 2024; Song et al., 2024) generate objects with ID (identity) preserved in the target scene, but they fall short in processing complicated 3D geometry relations (e.g., the occlusion) as they only consider 2D-level composition. To sum up, previous methods mainly either 1) fail to achieve both ID preservation and background harmony, or 2) do not explicitly take into account the geometry behind the background and fail to accurately composite objects and backgrounds in complex spatial relations.
@@
We divide the training procedure into two stages. In the first stage, we finetune an MLLM (e.g., LLaVA (Liu et al., 2023a, 2024)) for 2.5D predictions of objects in complex scenes with language instructions. In the second stage, we train the image composition model. To composite images with complicated spatial relationships, we introduce depth maps as conditions for image generation. In addition, we leverage an ID extractor to generate discriminative ID tokens and a frequency-aware detail extractor to extract the high-frequency 2D shape information (detail maps) for ID preservation. We unify the depth maps, ID tokens, and detail maps to guide a pre-trained text-to-image diffusion model to generate desired image composition results. This two-stage training paradigm allows us to utilize a number of existing 2D datasets for common visual tasks e.g., visual segmentation, and detection, avoiding collecting large-volume text-image data specially designed for arbitrary object-level image compositing.
@@ 2.1 Image compositing with Diffusion Models
Image compositing, an essential technique in image editing, seamlessly integrates a reference object into a background image, aiming for realism and high fidelity. Traditional methods, such as image harmonization (Jiang et al., 2021; Xue et al., 2022; Guerreiro et al., 2023; Ke et al., 2022) and blending (Pérez et al., 2003; Zhang et al., 2020, 2021; Wu et al., 2019) primarily ensure color and lighting consistency but inadequately address geometric discrepancies. The introduction of diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015; Song et al., 2021; Rombach et al., 2022) has shifted focus towards comprehensive frameworks that address all facets of image compositing. Methods like (Yang et al., 2023; Song et al., 2023) often use CLIP-based adapters to utilize pretrained models, yet they compromise the object’s identity preservation, focusing mainly on high-level semantic representations.More recent studies prioritize maintaining the appearance in generative object compositing. Notable developments in this field include AnyDoor (Chen et al., 2024) and ControlCom (Zhang et al., 2023a). AnyDoor integrates DINOv2 (Oquab et al., 2023) with a high-frequency filter, while ControlCom introduces a local enhancement module, both improving appearance retention. However, these approaches still face challenges in spatial correction capabilities. Most recent work IMPRINT (Song et al., 2024) trains an ID-preserving encoder that enhances the visual consistency of the object while maintaining geometry and color harmonization. However, none of the work can deal with composition with occlusion and more complex spatial relations. In contrast, our models novelly propose a 3D-aware generative model that allows more accuracy and complex image compositing while maintaining geometry and color harmonization.
@@ 3 Method
Figure 5: Overview of training pipeline of Bifröst on image compositing stage. A segmentation module is first adopted to get the masked image and object without background, followed by an ID extractor to obtain its identity information. The high-frequency filter is then applied to extract the detail of the object, stitch the result with the scene at the predicted location, and employ a detail extractor to complement the ID extractor with texture details. We then use a depth predictor to estimate the depth of the image and apply a depth extractor to capture the spatial information of the scene. Finally, the ID tokens, detail maps, and depth maps are integrated into a pre-trained diffusion model, enabling the target object to seamlessly blend with its surroundings while preserving complex spatial relationships.
@@
where $\mathcal{E}_{d}$ is the depth extractor and $\boldsymbol{c}_{\mathrm{d}}$ is the extracted depth condition. In our experiment, the detail and depth extractors utilize ControlNet-style UNet encoders, generating hierarchical detail maps at multiple resolutions.
@@
$\left.\mathcal{L}_{composite}=\mathbb{E}_{\boldsymbol{z}_{0},\boldsymbol{t},\boldsymbol{c}_{i},\boldsymbol{c}_{\mathrm{f}},\epsilon\sim\mathcal{N}(0,1)}\left[\|\epsilon-\epsilon_{\theta}\left(\boldsymbol{z}_{t},\boldsymbol{t},\boldsymbol{c}_{i},\boldsymbol{c}_{\mathrm{f}}\right)\right)\|_{2}^{2}\right],$ (5)
where $\boldsymbol{c}_{f}=\boldsymbol{c}_{h}+\lambda\cdot\boldsymbol{c}_{d}$ and $\lambda$ is a hyper parameter weight between two controls. Specifically, the ID tokens are injected into each UNet layer via cross-attention. The high-frequency detail and depth maps are first added together and then concatenated with the UNet decoder features at each resolution. During training, the pre-trained UNet encoder parameters are frozen to retain priors, while the decoder is fine-tuned to adapt to the new task.
@@
Dataset Generation. The ideal training data consists of image pairs capturing the same object in different scenes and poses, which existing datasets lack. Previous works (Yang et al., 2023; Song et al., 2023) use single images with augmentations like rotation and flip, which are insufficient for realistic pose and view variants. To address this, we use video datasets to capture frames of the same object as complementary following (Chen et al., 2024; Song et al., 2024). As illustrated in Fig. 6, our data preparation pipeline selects two frames from a video and extracts foreground masks. One frame is masked and cropped around the object to serve as the reference, while the other frame—with an augmented mask shape—acts as the background. The unmasked version of this second frame serves as the training ground truth. The dataset quality is another key to better identity preservation and pose variation. The full data used is listed in Tab. 1, which covers a large variety of domains such as nature scenes (SAM (Kirillov et al., 2023), LVIS (Gupta et al., 2019), HFlickr (Cong et al., 2020), DUT (Wang et al., 2017), and MSRA-10K (Borji et al., 2015)), panoptic video segmentation datasets (YoutubeVOS (Xu et al., 2018), VIPSeg (Miao et al., 2022), and MOSE (Ding et al., 2023)), and virtual try-on dataset (VitonHD (Choi et al., 2021)).
Figure 6: Data preparation pipeline of leveraging videos. Given a clip, we first sample two frames, selecting an instance from one frame as the reference object and using the corresponding instance from the other frame as the training supervision.
@@ 4 Experiment
Table 4: User study on the comparison between our Bifröst and existing alternatives. “Quality”, “Fidelity”, “Diversity”, and “3D Awareness” measure synthesis quality, object identity preservation, object local variation (i.e., across four proposals), and spatial relation awareness (i.e., occlusion) respectively. Each metric is rated from 1 (worst) to 5 (best).
@@ 4.1 Implementation Details
Zoom-in Strategy. During inference, given a scene image and a location box, we expand the box into a square using an amplification ratio of 2.0. The square is then cropped and resized to $512\times 512$ for input into the diffusion model. This approach enables handling scenes with arbitrary aspect ratios and location boxes covering extremely small or large areas.
@@
Evaluation metrics. We test the IoU and MSE loss to evaluate the accuracy of the predicted bounding box and depth of our MLLM. For the image compositing model, we evaluate performance on our constructed DreamBooth dataset by following DreamBooth (Ruiz et al., 2023) to calculate the CLIP-Score and DINO-Score, which measure the similarity between the generated region and the target object. We also compute the FID (Heusel et al., 2017) to assess realism and compositing quality. Additionally, we conduct user studies with 30 annotators to rate the results based on fidelity, quality, diversity, and 3D awareness.
@@ 5 Conclusion
Acknowledgements This work is partially supported by the National Natural Science Foundation of China (Grant No. 62306261), CUHK Direct Grants (Grant No. 4055190), and The Shun Hing Institute of Advanced Engineering (SHIAE) No. 8115074. We thank Boqing Gong (Google) for feedback on an early draft of this paper and Yuanyuan Zhang (University of Liverpool) for proposing the name Bifröst of our model.
@@ Appendix A Implementation Details
For 3D-aware image compositing model. We use Stable Diffusion 2.1 as the pre-trained diffusion model. The model can also be initialized from the pre-trained weights of  (Chen et al., 2024) to better utilize the prior knowledge. We choose Adam (Kingma and Ba, 2014) optimizer with an initial learning rate of $1e^{-5}$. The DINO-V2 (Oquab et al., 2023) ID extractor and the encoder of the U-net are frozen during the training. The decoder of the U-net, detail extractor, and depth extractor are fine-tuned during this process. The batch size is set as 16, we train 20k steps on 4 A800 GPUs, which takes 1 day to finish the training.
@@ B.1 MLLM Fine Tuning
We employ pre-trained visual encoders to extract the target object’s identity. Rather than using CLIP (Radford et al., 2021) for object embedding (Yang et al., 2023; Song et al., 2023), prior work (Chen et al., 2024; Song et al., 2024) demonstrates that DINO-V2 (Oquab et al., 2023) better captures discriminative features, projecting objects into an augmentation-invariant feature space. Thus, we use DINO-V2 to encode the image into a global token $\mathbf{T}{\text{g}}^{1\times 1536}$ and patch tokens $\mathbf{T}{\text{p}}^{256\times 1536}$. These tokens are concatenated to retain more information. Following (Chen et al., 2024), a linear layer bridges the DINO-V2 and pre-trained text-to-image UNet embedding spaces, yielding final ID tokens $\mathbf{T}_{\text{ID}}^{257\times 1024}$.
Collage Representation. Unlike prior methods (Casanova et al., 2023; Sarukkai et al., 2024; Chen et al., 2024) that use full-color images or high-frequency maps as complementary information, we found that these approaches often produce objects that resemble the reference image too closely, creating a "copy-paste" effect. To address this, we remove color information from the HF-map in (Chen et al., 2024), enhancing fine-grained details and improving the visual coherence of the generated objects.
Mask Shape Augmentation. To provide greater user control, we define five levels of coarse masks, including a bounding box mask, as shown in Fig. 14. As the mask level increases (from 1 to 5), the model gains more flexibility in object generation.
@@
The Detail Extract process is formulated as:
@@
where $\mathbf{K}_{h}$ and $\mathbf{K}_{v}$ are horizontal and vertical Sobel kernels (Kanopoulos et al., 1988) serving as high-pass filters. Here, $\otimes$ and $\odot$ represent convolution and Hadamard product, respectively. Given an object image $\mathbf{I}_{obj}$, high-frequency regions are extracted through these filters, and a shape-augmented mask $\mathbf{M}_{\text{aug}}$ is applied to remove information near the outer contour of the object.